a charge q is disturbed over 2 hallow spheres of raddi r and R(R>r) such that surface densites are equal fin potential at the common centre

Grade:12th pass

1 Answers

Khimraj
6 years ago
Let the charge on sphere of radius r is q1 then charge on sphere of radius R is q2 = q-q1.
Acccording to question surface charge density of both sphere are equal.
So (q1/4πr2) = (q-q1)/4πR2
Solving above equation q1 = q(r2)/(r2+R2)
and q2 = q – q1 = q(R2)/(r2+R2)
Potential at there common centre V = Kq1/r + Kq2/R
putting value of q1 and q2
V = Kq(r +R)/(r2+R2).
Hope it clears. If still have doubt you can ask me. If you like answer then  please approve it.